How much acid at 70% concentration must be added to 12 ounces at 40% concentration in order to obtain a mixture that is 50% acid?
-I know the answer will be closer to 40 than 70, but I'm just not sure how to work the problem out.

Answer by ptaylor(2198)

YOU R RIGHT ON!!! IT'S A LOT CLOSER TO 4 THAN TO 7.
Let x=amt of 70% needed
Then 12-x=amt of 40% needed
Now we know that the amt of pure acid before the mixture takes place has to equal the amt of pure acid after the mixture takes place, sooooo:
0.70x+0.40(12-x)=0.50*12
0.70x+4.8-0.40x=6
0.30x=1.2
x=4 oz------------ amt of 70% needed
12-x=12-4=8 oz------------amt of 40% needed
CK
0.70*4+0.40*8=0.50*12
2.8+3.2=6
6=6
Hope this helps---ptaylor